OpenStudy (anonymous):

what are the steps to get to my answer to this problem...25% of 2000

OpenStudy (anonymous):

25 % of 2000 =

OpenStudy (anonymous):

wt next u will write

OpenStudy (anonymous):

that is u have to find out what is equal to 25 % of 2000 is i am right GT Bro

OpenStudy (anonymous):

Yep! :-)

OpenStudy (anonymous):

umm confused

OpenStudy (agreene):

0.25*2000=500

OpenStudy (anonymous):

x% of something = x/100 times that thing. So, 25% of 2000 = 25/100 times 2000. Rest is simple arithmetic.

OpenStudy (anonymous):

thanks GT buddy